A bar measures 12 mm x 20 mm x 1 m. It has a specific gravity of 2.78. Determine its mass


1
Expert's answer
2021-10-31T18:16:23-0400

Given:

ρ=2.78=2780kg/m3\rho=2.78=2780\:\rm kg/m^3

V=12mm×20mm×1m=2.4104m3V=\rm 12\: mm\times20\: mm\times 1\: m=2.4*10^{-4}\: m^3


The mass of a bar

m=ρV=2780kg/m32.4104m3=0.667kgm=\rho V=\rm 2780\: kg/m^3*2.4*10^{-4}\: m^3=0.667 \: kg